社会医学 · 疫学 · 医学統計概論 2 単位 (選択) 1 年 (後期)

Introduction to social medicine, epidemiology and biostatistics  有澤孝吉 · 教授/ 大学院ヘルスバイオサイエンス研究部

伊藤博夫 · 教授/ 大学院ヘルスバイオサイエンス研究部

，上野修一 · 教授

，上村浩一 · 講師/ 大学院ヘルスバイオサイエンス研究部

，山内あい子 · 教授/ 医療生命薬学専攻

谷岡哲也 · 教授/ 大学院ヘルスバイオサイエンス研究部， �村彰 · 教授/ 医療生命薬学専攻

，日野出大輔 · 教授/ 大学院ヘルスバイオサイエンス研究部

森口博基 · 教授/ 大学院ヘルスバイオサイエンス研究部，楊河宏章 · 准教授/ 病院

【授業目的】社会医学・薬学，歯学等に関して，以下のような諸問題を多方面から概説する． 1. 医療の効果の科学的な分析において必要な医学統計学の基本について概説する． 2. 医療における IT の役割と病院マネジメントについて概説する． 3. 精神保健福祉に関する事象を明らかにするための分析手法を概説する． 4. 口腔の健康を保持増進し，口腔疾患を予防するために必要な疫学・統計の知識を概説する． 5. 医薬品の情報管理と安全対策はどうあるべきかを理解し，その具対策を概説する． 6. 薬害とは何か，どうして発生したか，どうすれば防止できるか等について概説する． 7. 論文作成のための調査デザイン，統計解析について概説する．

【授業概要】 1) 臨床試験の意義と実際について，科学的なデザイン，倫理的に必要な事項，日本での臨床試験の現状などの面から概説する． 2) 大学病院における IT の活用事例や病院経営におけるマネジメントの方法について概説する． 3) 精神保健福祉における事象の特徴を統計学的に明らかにするための方法論について，先行研究も踏まえて説明する． 4) 口腔疾患について，発生要因の追求や疾病対策の効果判定に関する事例をまじえながら教授し，科学的根拠に基づいた口腔疾患の予防法を概説する． 5) 患者の権利として要求される医療 (医薬品) 情報開示・提供に際しての情報管理の在り方や医薬品の安全対策とリスクマネージメントの具体論等を概説する． 6) 多くの薬害事例について，それらの発生要因や背景を考察し，薬害を如何に防止するかを考える．

7) データの種類および取り扱い，統計的検定の意味など，統計解析の基本について概説する．また， SPSS を用いたデータ解析の実習を行う．

【授業形式】 [形態] 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修上の注意】講義は e-learning 化しているので， e-learning 学習も出席として取

り扱う．【到達目標】社会医学 · 薬学 · 歯学等に関して，授業目的に示した講義内容の理解

が深まることを目標にする．【授業計画】

1. 論文作成のためのデータの基本的な取り扱いと統計解析について (担当者: 上村浩一)

2. 論文作成のためのデータの基本的な取り扱いと統計解析について (担当者: 上村浩一)

3. 学位論文作成のための疫学，統計解析の実際 (SPSS) (担当者: 有澤孝吉)

4. 学位論文作成のための疫学，統計解析の実際 (SPSS) (担当者: 有澤孝吉)

5. 臨床試験の意義と実際について (担当者: 楊河宏章)

6. 臨床試験の意義と実際について (担当者: 楊河宏章)

7. 病院における IT の役割 I (担当者: 森口博基)

8. 病院における IT の役割 II (担当者: 森口博基)

9. 精神保健福祉に関する問題の統計学的分析手法 (担当者: 上野修一)

10. 精神保健福祉に関する問題の統計学的分析手法 (担当者: 谷岡哲也)

11. 齲蝕・歯周病の疫学，口腔疾患と全身の健康 (担当者: 伊藤博夫)

12. 齲蝕・歯周病の疫学，口腔疾患と全身の健康 (担当者: 日野出大輔)

13. 医薬品の情報管理，医薬品の安全対策 (担当者: 木原勝)

14. 医薬品の情報管理，医薬品の安全対策 (担当者: 木原勝)

15. 薬害とは何か，なぜ起こるのか，どうすれば薬害を防げるか (担当者: 徳村彰)

16. 薬害とは何か，なぜ起こるのか，どうすれば薬害を防げるか (担当者: 徳

村彰) 【成績評価】講義中に提示した課題についてレポートを提出する．各講師の先生

の出席点・評点を総合的に考慮して決定する．【再試験】再試験は基本的には行わない．【教科書】各講師の指定した教材を使用する．【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218231 【連絡先】

⇒ 有澤 (088-633-7071, arisawa@basic.med.tokushima-u.ac.jp) ���� (オフィスアワー: e-mailにてアポイントメントを取り，面談してください． )

【備考】講師の先生によって講義をする場所が異なるので注意すること．

Introduction to social medicine, epidemiology and biostatistics 2 units (selection) 1st-year(2nd semester)

Kokichi Arisawa · Professor / Institute of Health Biosciences , Hiro-O Ito · Professor / Institute of Health Biosciences , Shu-ichi Ueno · Professor

, Hirokazu Uemura · Associate Professor / Institute of Health Biosciences

Aiko Yamauchi · Professor / Pharmaceutical Life Sciences , Tetsuya Tanioka · Professor / Institute of Health Biosciences

, Akira Tokumura · Professor / Pharmaceutical Life Sciences

Daisuke Hinode · Professor / Institute of Health Biosciences , Hiroki Moriguchi · Professor / Institute of Health Biosciences

, Hiroaki Yanagawa · Associate Professor / University Hospital

Target⟩ 1, Clinical trial as a tool to establish clinical evidence 2. The role of the IT in the medical treatment and the hospital management are outlined. 3. Outline analytical skills for clarifying the phenomenon about Mental Health and Welfare. 4. Outline epidemiology and analytical methods for oral health promotion and/or oral disease prevention. 6. Outline damages from medicines and its etiology and prevention. 7. Outline designing of surveys and statistical analysis of data for writing a thesis.

Outline⟩ 1) Principles and practice of clinical trial from scientific and ethical viewpoint will be reviewed. 2) The usage of IT and the method of the management in the university hospital are outlined. 3) We will explain the statistical method for clarifying the feature of phenomenon based on previous work in Mental Health and Welfare. 4) We will explain the etiology of oral disease and its prevention based on the evidence of oral sciences. 6) Instruct background and etiology of several examples of damages from medicines and discuss possible prevention of the damages. 7) Basic aspects of statistical analysis, such as type of data and appropriate treatment of data, and meaning of hypothesis testing, will be summarized. In addition, practice of data analysis will be done using SPSS.

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Notice⟩ It is possible to take a lesson by e-learning. Goal⟩ The goal is to deepen one’s understanding of contents presented during

lectures, regarding social medicine, pharmaceutical science and dentistry.

Schedule⟩

1. Basic handling and statistical analysis of data for article making (担当者: Hirokazu Uemura)

2. Basic handling and statistical analysis of data for article making (担当者: Hirokazu Uemura)

3. Practice of epidemiology and statistical data analysis for writing a thesis (担当者: Kokichi Arisawa)

4. Practice of epidemiology and statistical data analysis for writing a thesis (担当者: Kokichi Arisawa)

5. Principles and practice of clinical trial (担当者: Hiroaki Yanagawa)

6. Principles and practice of clinical trial (担当者: Hiroaki Yanagawa)

7. The role of IT in a hospital (担当者: Hiroki Moriguchi)

8. The role of IT in a hospital (担当者: Hiroki Moriguchi)

9. Statistical procedure for analyzing issues facing the Mental Health and Welfare (1) (担当者: Shu-ichi Ueno)

10. Statistical procedure for analyzing issues facing the Mental Health and Welfare (2) (担当者: Tetsuya Tanioka)

11. Epidemiology of dental caries and periodontal disease, and the relationship between oral disease and systemic disease (担当者: Hiro-o Ito)

12. Epidemiology of dental caries and periodontal disease, and the relationship between oral disease and systemic disease (担当者: Daisuke Hinode)

13. 医薬品の情報管理，医薬品の安全対策 (担当者: 木原勝)

14. 医薬品の情報管理，医薬品の安全対策 (担当者: 木原勝)

15. What is damages from medicines? Why did the damages occurred? How to prevent the damages? (担当者: Akira Tokumura)

16. What is the damages from medicines? Why did the damages occurred?

How to prevent the damages? (担当者: Akira Tokumura) Evaluation Criteria⟩ Reports should be submitted on questions presented during lectures. Grades are determined by summing points from all instructors.

Re-evaluation⟩ In principle, there is no reexaminations.

Textbook⟩ Teaching materials designated by each instructor wil be used.

Reference⟩ [参考資料]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218231 Contact⟩

⇒ Arisawa (+81-88-633-7071, arisawa@basic.med.tokushima-u.ac.jp) ����

(Office Hour: Please contact each instructor by an e-mail.)

Note⟩ Different classrooms may be used according instructors.